During the month of February the following transactions occurred to produce 3 products.

1. Material purchase on account 

Sand 500.00

Steel 790.00

Chemical 4500.00

2. Direct material issued and direct labor cost 

  Direct material issued direct labor cost

Job – 1. 227.50. 160.00

Job – 2. 780.00. 210.00

Job – 3. 3900.00. 175.00

3. Factor over head is 60% of direct labor cost 

4. Finished goods sold Birr 7200.00

Required 

1. Record journal entries?

2. Calculate cost of each job ?

3. Calculate gross profit ?


1
Expert's answer
2021-03-29T12:19:48-0400

1.debit materials credit payments to suppliers- 500

debit materials credit payments to suppliers- 790

debit materials credit payments to suppliers- 4500

debit main production credit materials - 227.50

debit main production credit materials - 780

debit main production credit materials - 3900

debit main production credit settlements with personnel - 160

debit main production credit settlements with personnel - 210

debit main production credit settlements with personnel - 175

debit main production credit general expenses - 96 (160*0.6)

debit main production credit general expenses - 126 (210*0.6)

debit main production credit general expenses - 105 (175*0.6)

debit payments to customers credit sales- 7200

2.

Job – 1: 227.50+160+96=483.50

Job – 2: 780+210+126=1 026

Job – 3: 3900+175+105=4 180

3.gross profit=Revenue-costs=7200-(483.50+1026+4180)=1 510.5